OpenStudy (anonymous):

Let y=e^(x/5). Find the differential dy when x=3 and dx=0.3 Find the differential dy when x=3 and dx=0.01

OpenStudy (anonymous):

y=e^x/5 dy/dx= 1/5 e^(x/5)

OpenStudy (anonymous):

\[\color{green}{\text{hello math girl!}}\]

OpenStudy (anonymous):

hey! i tried to im u

OpenStudy (anonymous):

dy= 1/5 e^(x/5) dx now plug them in